# Dynamic Avalanches: Rate-Controlled Switching and Race Conditions

### v1

@@ Acknowledgements.
We thank D. Shohat, P. Baconnier, J. Liu, and M. Liu for insightful discussions, and D. Ursem for for technical support. L. J. and M. v. H. acknowledge funding from European Research Council Grant ERC-101019474.
@@ Appendix A Experimental Methods and Samples
The mechanical response of our samples is characterized using a horizontal tensile testing machine (Instron 5965), which provides precise control over axial displacement U with an accuracy better than 4 $\mu$m. A 100 N load cell is employed to measure the applied force with a resolution down to 0.5 mN. All experiments are conducted at a constant displacement rate of 10 mm/min. During testing, the samples are imaged using both a Canon camera (EOS 800D) and a high-speed camera (Phantom v4.2) to capture their dynamic evolution. The samples are tested either in air or fully or partially immersed in Polydimethylsiloxane oil (viscosity: 350 cSt) for various damping conditions. To implement partial submersion of a hysteron, we mount two wings (green parts in Fig. 3c) to its sides. The wings are then immersed in silicone oil while the rest of the sample remains clear of the oil, creating a partially damped environment for the sample.
@@ Appendix B Numerical Model
To investigate the dynamical behavior of coupled bistable elements, we perform simulations of a simple mechanical model introduced recently [27, 28]. This model represents networks of bistable springs, linear springs, masses and viscous dampers (Fig. 1d). First, each bistable spring is represented with a cubic force-displacement curve $f_{i}(u_{i})$. Second, the elliptical spring in our samples is represented by a linear spring $f_{k}(u_{k})$. The detailed forms of $f_{i}(u_{i})$ and $f_{k}(u_{k})$ can be obtained from fitting to finite element simulations of our samples, or can be adjusted to represent the range of stability of the states [27, 28, 16, 26]. Together, networks of these elements describe the stable and unstable collective states and exhibit hysteretic transitions between stable states.
@@ Appendix C Design of Three-Bit Counter
We consider the design of a targeted pathway that materializes a three-bit binary counter (Fig. A1a). We start from a simple model of three serially coupled bilinear elements, each describe by [16, 26]:
@@
$\dots$<⋯ < $f_{3}^{-}\!+\!g_{2}\!+\!g_{3}<f_{3}^{+}\!+\!g_{2}<f_{1}^{-}\!+\!g_{1}$$\dots$
$\dots$<⋯ < $f_{1}^{+}\!+\!g_{2}\!+\!g_{3}<f_{3}^{-}\!+\!g_{1}\!+\!g_{3}<f_{3}^{+}\!+\!g_{1}$$\dots$
$\dots$<⋯ < $f_{2}^{-}\!+\!g_{1}\!+\!g_{2}<f_{2}^{+}\!+\!g_{1}\!+\!g_{3}$$\dots$
$\dots$<⋯ < $f_{3}^{-}\!+\!g_{1}\!+\!g_{2}\!+\!g_{3}<f_{3}^{+}\!+\!g_{1}\!+\!g_{2}~{}.$ (5)
There is a range of choices for the switching forces and force drops $g_{i}$—which both can be designed experimentally at the level of individual elements—that satisfy these equations. For example, an equidistant ordering of the switching fields requires rapidly growing values of the force drop $g_{i}$ with significance of the elements. Hence, in our experimental design we take each element to satisfy the nested constraints on the critical forces, and the rapidly increasing values of the force drops. After some iterative design, this produced sample 3, which satisfies both sets of design equations (Fig. A1c). Note that while sample 3 exhibits sharp transitions between states in the correct order, the deformation of the largest element exhibits higher beam modes, which modify $f_{1}(u_{1})$ and make the collective curve $F(U)$ deviate from the strictly piecewise linear target - we stress that the sequence of transitions is correct. This deviation shows the robustness of our design strategy to such experimental deviations.

### v3

@@ Acknowledgements.
We thank D. Shohat, P. Baconnier, J. Liu, and M. Liu for insightful discussions, and D. Ursem for technical support. L. J. and M. v. H. acknowledge funding from the European Research Council Grant ERC-101019474.
@@ Appendix A Experimental Methods and Samples
[…]
@@ Appendix B Numerical Model
To investigate the dynamical behavior of coupled bistable elements, we perform simulations of a simple mechanical model introduced recently [30, 31]. This model represents networks of bistable springs, linear springs, masses and viscous dampers (Fig. 1d). First, each bistable spring is represented with a cubic force-displacement curve $f_{i}(u_{i})$. Second, the elliptical spring in our samples is represented by a linear spring $f_{k}(u_{k})$. The detailed forms of $f_{i}(u_{i})$ and $f_{k}(u_{k})$ can be obtained from experimental data of our samples, or can be adjusted to represent the range of stability of the states [30, 31, 16, 29]. Together, networks of these elements describe the stable and unstable collective states and exhibit hysteretic transitions between stable states.
@@ Appendix C Design of Three-Bit Counter
We consider the design of a targeted pathway that materializes a three-bit binary counter (Fig. A1a). We start from a simple model of three serially coupled bilinear elements, each described by [16, 29]:
@@
$\dots<$ $f_{3}^{-}\!+\!g_{2}\!+\!g_{3}<f_{3}^{+}\!+\!g_{2}<f_{1}^{-}\!+\!g_{1}$ $\dots$
$\dots<$ $f_{1}^{+}\!+\!g_{2}\!+\!g_{3}<f_{3}^{-}\!+\!g_{1}\!+\!g_{3}<f_{3}^{+}\!+\!g_{1}$ $\dots$
$\dots<$ $f_{2}^{-}\!+\!g_{1}\!+\!g_{2}<f_{2}^{+}\!+\!g_{1}\!+\!g_{3}$ $\dots$
$\dots<$ $f_{3}^{-}\!+\!g_{1}\!+\!g_{2}\!+\!g_{3}<f_{3}^{+}\!+\!g_{1}\!+\!g_{2}~.$ (5)
There is a range of choices for the switching forces and force drops $g_{i}$—which both can be designed experimentally at the level of individual elements—that satisfy these equations. For example, an equidistant ordering of the switching fields requires rapidly growing values of the force drop $g_{i}$ with the significance of the elements. Hence, in our experimental design we take each element to satisfy the nested constraints on the critical forces, and the rapidly increasing values of the force drops. After some iterative design, this produced sample 3, which satisfies both sets of design equations (Fig. A1c). Note that while sample 3 exhibits sharp transitions between states in the correct order, the deformation of the largest element exhibits higher beam modes, which modify $f_{1}(u_{1})$ and make the collective curve $F(U)$ deviate from the strictly piecewise linear target - we stress that the sequence of transitions is correct. This deviation shows the robustness of our design strategy to such experimental deviations.